OpenStudy (anonymous):

Solve for the variable in the equation 16 = t + 7 A: 23 B: 12 C: 13 D: 9

OpenStudy (anonymous):

D: 9

OpenStudy (anonymous):

16 = t + 7 Subtract 7 from both sides. 16 - 7 = t +7 - 7 16-7 = t 9 = t